An archeologist found the remains of an ancient wheel, which she then placed on a grid. If an arc of the wheel passes through A(-7, 0), B(-3, 4) and C(7, 0) locate the center of the wheel, and the standard equation of the circle defining its boundary.


1
Expert's answer
2020-12-03T08:00:21-0500

the circle has an equation "(x-a)^2+(y-b)^2 = R^2" ,

(a, b) - center circle, point A, B, C belong circle

"\begin{cases}
(-7-a)^2+(0-b)^2 = R^2\\
(7-a)^2+(0-b)^2 = R^2\\
(-3-a)^2+(4-b)^2 = R^2\\
\end{cases}\\
\begin{cases}
49+14a+a^2+b^2 = R^2\\
49-14a+a^2+b^2 = R^2\\
(-3-a)^2+(4-b)^2 = R^2\\
\end{cases}\\
\begin{cases}
28a = 0\\
49-14a+a^2+b^2 = R^2\\
(-3-a)^2+(4-b)^2 = R^2\\
\end{cases}\\
\begin{cases}
a = 0\\
49-14a+a^2+b^2 = R^2\\
(-3-a)^2+(4-b)^2 = R^2\\
\end{cases}\\
\begin{cases}
49+b^2 = R^2\\
9+(4-b)^2 = R^2\\
\end{cases}\\
\begin{cases}
49+b^2 = R^2\\
9+16-8b+b^2 = R^2\\
\end{cases}\\
\begin{cases}
49+b^2 = R^2\\
25-8b+b^2 = R^2\\
\end{cases}\\
24+8b = 0\\
b = -3\\
R^2 = 58\\
x^2 +(y+3)^2 = 58"

answer: (0, -3)

"x^2 +(y+3)^2 = 58"